OpenStudy (anonymous):

1/4 divided by 9/4 = ? 1 7/9 9 9/16 1/9

OpenStudy (anonymous):

(1/4) / (9/4) =(1/4)(4/9) =1/9

OpenStudy (anonymous):

thanks :)